# Differentially Private Federated Learning: A Systematic Review

### v3

@@ 2.1.2. Vertical Federated Learning (VFL)
Compared to HFL which processes data from multiple clients sharing the same feature set, VFL trains a better predictor by combining all features from different clients. VFL assumes that the data is partitioned across feature space. The data vector $x_{i}\in\mathbb{R}^{1\times d}$ is distributed among $K$ participants $\{x_{i,k}\in\mathbb{R}^{1\times d_{k}}\}_{k=1}^{K}$ in $D=\{x_{i},y_{i}\}_{i=1}^{N}$, where $d_{k}$ is the feature dimension of the $k^{th}$ participant, $k\in[K-1]$, and the $K^{th}$ participant has the label information $y_{i}=y_{i,K}$ (Yang et al., 2019). The participant with the label, the $K^{th}$ participant, is referred to as the server, while the others are referred to as clients 777 Some articles refer to them as “active” and “passive” parties (Liu et al., 2024); however, to maintain consistency with HFL, we refer to them as the “server” and “clients” here.. Each client $k$ has a dataset $D_{k}\triangleq\{x_{i,k}\}_{i=1}^{N}$, and the server has a dataset $D_{K}\triangleq\{x_{i,K},y_{i,K}\}_{i=1}^{N}$.
The global model $\mathbf{w}$ decomposes into the client models $\mathbf{w}_{k}$, $k\in[K-1]$, and the server model $\mathbf{w}_{K}$. Each $\mathbf{w}_{k}$ accesses dataset $D_{k}$ for forward propagation to yield the result $h_{k}$, while $\mathbf{w}_{K}$ accesses $D_{K}$ for both forward and backward propagation. The objective of VFL is to collaboratively train a global model $\mathbf{w}^{\star}$ with all clients, aiming to minimize the global objective function $F\left(\cdot\right)$, essentially seeking to find:
(2) $\mathbf{w}^{\star}$ $\triangleq\min_{\mathbf{w}}F\left(\mathbf{w}\right),\,\text{where}\,F\left(\mathbf{w}\right)\triangleq\frac{1}{N}\sum_{i=1}^{N}L_{K}(\mathbf{w}_{K};h_{1}\left(x_{i,1},\mathbf{w}_{1}\right),\ldots,h_{K}\left(x_{i,K},\mathbf{w}_{K}\right),y_{i,K}),$
where $L_{K}(\cdot)$ donates the loss function of the server.
@@ 2.1.3. Transfer Federated Learning (TFL)
TFL was proposed by Liu et al. (Liu et al., 2020c), it refers to the scenario in FL settings where there is limited or no overlap in the feature space or sample space between clients. In this scenario, there are often multiple source domain datasets $D_{k}\triangleq\left\{\left(x_{i,k},y_{i,k}\right)\right\}_{i=1}^{N_{k}},k\in[K-1]$, and a target domain dataset $D_{K}\triangleq\left\{x_{i,K}\right\}_{i=1}^{N_{K}}$. Here, $D_{k}$ and $D_{K}$ are held by different participants, with those holding the source domain datasets called clients and those holding the target domain dataset called the server. Under this setup, the goal is to enable multiple parties to jointly train a model that can generalize to the target domain dataset held by the server, without exposing data to each other. In most cases, there are often few or even no labels in target domain dataset, in which case it need to use the source domain datasets from each client to generate pseudo-labels for the target domain dataset firstly.
@@ 2.2. DP Models
Differential privacy (Dwork et al., 2006) is a rigorous mathematical framework that formally defines data privacy, originally used for privacy protection techniques on data collection servers. It can resist differential attacks using statistical queries by adding an appropriate amount of noise to the output of statistical queries, making it almost impossible for adversaries to discern the statistical differences between two adjacent datasets. There is a trusted server in DP setting. It requires that a single entry in the input dataset must not lead to statistically significant changes in the output if differential privacy holds (Dwork et al., 2006, 2014). The formal definition of DP is presented as follows.
@@ 2.2.2. LDP
From Definition 2.4, it can be seen that LDP ensures that algorithm $\mathcal{R}$ satisfies $(\epsilon,\delta)$-LDP by controlling the similarity of any two records’ output. In short, according to a certain output result of privacy algorithm $\mathcal{A}$, it is almost impossible to infer which record its input data is. In DP, its privacy algorithm $\mathcal{A}$ is defined by neighbor dataset, so it requires truthful third-party data collectors to protect data analysis results. We can see that the main difference between DP and LDP lies in the definition of neighboring datasets.
LDP requires that all pairs of sensitive data must satisfy the same $\epsilon$ privacy guarantee, which may hide too much information from the datasets, leading to insufficient utility for certain applications. Gursoy et al. (Gursoy et al., 2019) extended metric-based extensions of differetial privacy to LDP and proposed condensed LDP (CLDP) to improve the utility, which measures the level of privacy guarantee between any pair of sensitive data based on their distance.
@@ 2.2.3. Shuffle Model
In recent years, research has introduced the shuffle differential privacy model, built upon the foundation of LDP. It adds a trusted shuffler [^8] Some researchers also assume the shuffler is semi-honest (Goldreich, 2009), where the cryptography-based protocol is required to perform a secure shuffle without accessing the values. Since there exist many secure shuffle protocols handling this case, we can simply assume it’s honest and pay attention on the part of DP. between the server and the clients, which shuffles the data items submitted by clients to achieve anonymization. In shuffle model, each client satisfies privacy guarantee of LDP when facing shuffler, and then achieves privacy guarantee of DP when facing the server since the shuffled data forms a dataset and generates the notion of neighborhood dataset.
@@ 2.2.4. Relation between DP, LDP and shuffle model
2.2.5. Basic Properties and Loss Composition Mechanisms
There are some basic properties that apply in both DP and LDP. The first one is post-processing, which means that any post-processing on the output of an $(\epsilon,\delta)$-DP algorithm will remain $(\epsilon,\delta)$-DP.
Lemma 2.6 (Post-processing (Dwork et al., 2014)).
Let $\mathcal{A}:\mathbb{X}\rightarrow\mathbb{Y}$ be a mechanism satisfying $(\epsilon,\delta)$-DP. Let $f:\mathbb{Y}\rightarrow\mathbb{R}$ be random function, then $f\circ\mathcal{A}:\mathbb{X}\rightarrow\mathbb{R}$ also satisfies $(\epsilon,\delta)$-DP.
Form the Lemma 2.6, we can see that privacy cannot be weakened by any post-processing. For example in FL, we can let the gradients satisfy the $(\epsilon,\delta)$-DP firstly, and because the gradient decent is a post-processing operation, the model parameters also satisfy the $(\epsilon,\delta)$-DP. The second property is the parallel composition property, that divides the dataset into disjoint chunks, and a differential privacy mechanism is run separately on each chunk. Since these chunks are disjoint, each individual’s data appears in only one chunk. Therefore, even if the differential privacy mechanism is executed multiple times, it only runs once on the data of each individual.
Lemma 2.7 (Parallel Composition Theorem (McSherry, 2009)).
Let mechanism $A$ consists of a sequence of adaptive mechanisms $\mathcal{A}_{1}(D_{1}),\mathcal{A}_{2}(D_{2}),...,\mathcal{A}_{k}(D_{k})$ that satisfy $\epsilon_{1}$-DP, $\epsilon_{2}$-DP, …, $\epsilon_{k}$-DP, respectively. And $D_{1},D_{2},...D_{k}$ are the disjoint chunks of dataset $D$. Then $A$ satisfies $(\max_{i\in[1,..,k]}\epsilon_{i},\max_{i\in[1,..,k]}\delta_{i})$-DP.
In differentilly private HFL, the dataset held by each client can be considered as an independent disjoint dataset, making it compatible with the principle of parallel composition. In differentilly private TFL, dividing privacy data into disjoint teacher datasets, such that the aggregation complies with the principle of parallel composition, is also a common practice. The third property is the sequential composition property, which allows the combination of privacy budgets from multiple privacy algorithms to satisfy an overall privacy budget.
Lemma 2.8 (Basic Sequential Composition Theorem (Dwork et al., 2014)).
If mechanism $A$ consists of a sequence of adaptive mechanisms $\mathcal{A}_{1}(D),\mathcal{A}_{2}(D),...,\mathcal{A}_{k}(D)$, that satisfy $\epsilon_{1}$-DP, $\epsilon_{2}$-DP, …, $\epsilon_{k}$-DP, respectively. $A$ satisfies $(\sum_{i=1}^{k}\epsilon_{i},\sum_{i=1}^{k}\delta_{i})$-DP.
Definition 2.9.
[…]
Definition 2.10.
Definition 2.11.
Based on Rényi divergence, zCDP can be obtained as follows:
Definition 2.12.
(6) $D_{\alpha}(\mathcal{A}(D)||\mathcal{A}(D^{\prime}))\leq R\alpha.$
And the following Lemma 2.13 defines the standard form for converting $(\alpha,R)$-zCDP to ($\epsilon$, $\delta$)-DP.
Lemma 2.13.
(Conversion from zCDP to DP (Bun and Steinke, 2016b)). if a randomized mechanism $A:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-zCDP ,then it satisfies$(R+2\sqrt{R\log(1/\delta)},\delta)$-DP for any $10<\delta<$ 1.
Definition 2.14.
(7) $D_{\alpha}(\mathcal{A}(D)||\mathcal{A}(D^{\prime}))\leq R.$
And the following Lemma 2.15 defines the standard form for converting $(\alpha,R)$-RDP to ($\epsilon$, $\delta$)-DP.
Lemma 2.15.
(Conversion from RDP to DP (Balle et al., 2020)). if a randomized mechanism $A:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
Definition 2.16.
Definition 2.17.
(9) $T(\mathcal{A}(D),\mathcal{A}(D^{\prime}))\geq G_{\mu},$
And the following Lemma 2.18 defines the standard form for converting $\mu$-GDP to ($\epsilon$, $\delta$)-DP.
Lemma 2.18.
2.3. Fundamental Mechanisms for Differentially Private FL
Data Types DP LDP
2.3.1. Perturbation Mechanisms for DP in FL
Gaussian Mechanism (Dwork et al., 2014) is the most popular mechanism for DP in FL. Assume $f:x\to\mathbb{X}$ be a function related to a query. For any $x\in\mathbb{X}$, this mechanism adds noise $n\sim\mathcal{N}(0,\sigma^{2})$ to the $f(x)$ to ensure $(\epsilon,\delta)$-DP. Here $\mathcal{N}(0,\sigma^{2})$ denotes the noise sample from the Gaussian (normal) distribution with mean 00 and variance $\sigma^{2}$, where $\sigma^{2}=\frac{2\Delta f^{2}\log(1.25/\delta)}{\epsilon^{2}}$ and $\Delta f$ is the sensitivity of $f$.
Definition 2.19.
Skellam Mechanism (Agarwal et al., 2021) is a response to the issue of the Discrete Gaussian mechanism not being closed under summation. Let $\epsilon>0,\alpha>1$ and $f:x\to\mathbb{X}$ be a function related to a query. For any $x\in\mathbb{X}$, this mechanism adds noise $n\sim\mathrm{Sk}_{0,\mu}$ to $f(x)$ to satisfy $(\alpha,\frac{\alpha\Delta f^{2}}{2\mu}+\min\left(\frac{(2\alpha-1)\Delta f^{2}+6\Delta f}{4\mu^{2}},\frac{3\Delta f}{2\mu}\right))$-RDP. Here $\Delta f$ is the sensitivity of $f$ and $\mathrm{Sk}_{0,\mu}$ denotes sampling from the Skellam distribution with mean 00 and variance $\mu$ as follows.
Definition 2.20.
where $I_{v}(x)$ is the modified Bessel function of the first kind.
2.3.2. Perturbation Mechanisms for LDP in FL
RAPPOR (Erlingsson et al., 2014), also building on RR, incorporates a Bloom filter to facilitate discrete frequency estimation in large domains. Specifically, given a discrete domain $\{1,2,\ldots,d\}$, denoted as $[d]$, and $k$ hash functions =$\mathbb{H}=\{H_{1},H_{2},\dots,H_{k}\}$ that map values from $[d]$ to $[k]$. Each user first uses these $k$ hash functions to map their value $v$ into the Bloom filter with $m$ bits and then perturbs each bit in the Bloom filter using RR. To reduce collisions, where two values are hashed to the same set of indices, RAPPOR allocates users into several cohorts, assigning a unique group of hash functions to each cohort. For longitudinal privacy, which protects against attacks during multiple accesses, RAPPOR splits $\epsilon$ into two parts: the first for permanent perturbation via RR and the second for instantaneous perturbation based on the permanent perturbation.
@@ 3.1. DP-HFL
(Client-level DP (CL-DP) [^9] Some articles treat “user-level DP” as equivalent to “client- level DP” because they assume that each client has only one user (Cheng et al., 2022; Yang et al., 2023b). However, we provide a more general definition, where we consider the possibility of one or more users under a single client. Further discussion and details regarding the definition and functionality of user-level are addressed in Section 6.). Under CL-DP, two datasets $D$ and $D^{\prime}$ are neighbouring if they differ in a single client or device (either through addition or through removal).
@@ 3.1.1. SL-DP
However, how to overcome the challenge of data heterogeneity of DP-FL is the concern of most of the work. Huang et al. (Huang et al., 2020) proposed differential privacy convolutional neural network with adaptive gradient descent algorithm (DPAGD-CNN) to update the training parameters of each client. They selected the best learning rate from a candidate set based on model evaluation in each round of local DPSGD. Noble et al. (Noble et al., 2022) introduced DP-SCAFFLOD, an extension of the SCAFFLOD algorithm (Karimireddy et al., 2020), which incorporates differential privacy. It employs a control variable to limit model drift during local DPSGD, aligning them more closely with the global model direction. And Using advanced results from DP theory and optimization to establish the convergence of DP-SCAFFLOD for convex and non-convex objectives. Wei et al. (Wei et al., 2021) found that there is an optimal number of communication rounds in terms of convergence performance for a given privacy budget $\epsilon$, which has motivated them to adaptively allocate privacy budgets in each round. Fu et al. (Fu et al., 2022a) proposed the Adap DP-FL algorithm, which includes adaptive gradient clipping and adaptive noise scale reduction methods. In the gradient clipping step of DPSGD, gradients are clipped using adaptive thresholds to account for the heterogeneity of gradient magnitudes across clients and training rounds. Additionally, during noise addition, the noise scale gradually decreases (or the privacy budget gradually increases) as gradients converge across training rounds. Yang et al. (Yang et al., 2023c) start from the Non-IID data itself, by concurrently updating local data transformation layers during local model training, thereby reducing the additional heterogeneity introduced by DP, and consequently improving the utility of FL models for Non-IID data. Their method was also applied to the CL-DP setting.
@@
The allocation of privacy budget is also the focus of the study, as the privacy budget determines factors such as the number of iteration rounds and noise scale, it directly impacts the performance of the trained model. Ling et al. (Ling et al., 2023) investigated how to achieve better model performance under constraints of privacy budget and communication resources. They conducted convergence analysis of DP-HFL, derived the optimal number of local iterations before each aggregation. Liu et al. (Liu et al., 2021b) considered scenarios with heterogeneous privacy budgets and proposed the Projected Federated Averaging (PFA) algorithm. This algorithm utilizes the top singular subspace of model updates submitted by clients with higher privacy budgets to project them onto model updates from clients with lower privacy budgets. Furthermore, it reduces communication overhead by having clients with lower privacy budgets upload projected model updates instead of original model values. Zheng et al. (Zheng et al., 2021) integrated GDP privacy metric into DP-HFL, proposing a private federated learning framework called PriFedSync. While considering the cost of communication, Li et al. (Li et al., 2022c) proposed SoteriaFL, which is a unified framework for compressed private FL. They use shifted compression scheme (Horváth et al., 2023) to compress the perturbed parameters for efficient communication while maintaining high utility.
@@ 3.1.2. CL-DP
In addition to this, many studies have focused on reducing the impact of noise on uploaded model parameters. Chen et al. (Cheng et al., 2022) observed that using a small clipping threshold can decrease the injected noise volume. They reduced the norm of local updates by regularizing local models and making the local updates sparse. DP-FedSAM  (Shi et al., 2023) uses a SAM optimizer (Foret et al., 2020) to help model parameters escape saddle points and enhance the robustness of the model to noise, aiming to find more stable convergence points. Bietti et al. (Bietti et al., 2022) proposed PPSGD, which trains personalized local models and ensures global model privacy simultaneously. They utilized personalized models to enhance the global model’s performance. Similarly, Yang et al. (Yang et al., 2023b) dynamically preserved high-information model parameters locally from noise impact using layer-wise Fisher information. They also introduced an adaptive regularization strategy to impose differential constraints on model parameters uploaded to the server, enhancing robustness to clipping. Xu et al. (Xu et al., 2023a) took into account that the size of the softmax layer of model parameters is linearly related to the number of sample labels. By keeping the softmax layer local and not uploading it, the participation of more clients (i.e., more sample labels) in federated learning does not result in more noise injection. Triastcyn et al. (Triastcyn and Faltings, 2019) leveraged the assumption that data is distributed similarly across clients in HFL, which leads to their updates being highly consistent, and utilized BDP for privacy auditing to obtain tighter privacy bounds. Their assumption and method is not only used under the CL-DP, but also applied in SL-DP scenario.
CL-DP with Secure Aggregation (SA)  [^12] Many works refer to this framework as “distributed dp” (Kairouz et al., 2021a; Agarwal et al., 2021; Yang et al., 2023c).. As mentioned above, CL-DP often requires a secure server for aggregation. In order to eliminate the reliance on a trusted central authority, many recent studies have combined differential privacy techniques with secure aggregation (Bonawitz et al., 2017) to achieve CL-DP. As shown in Figure 4, after local training, each client adds a small amount of differential privacy noise to the model parameters, then encrypts and sends them to the server. The server performs aggregation, enabling the aggregation of noise levels from each client, thereby ensuring that the encrypted model satisfies CL-DP. However, secure aggregation relies on modular arithmetic and cannot be directly compatible with continuous noise mechanisms (such as Laplace mechanism and Gaussian mechanism) currently used to implement differential privacy. This has led researchers to begin designing discrete noise addition mechanisms and integrating them with secure aggregation to characterize differential privacy.
@@
However, the above methods rely on per-parameter quantization, resulting in significant communication overheads. To further explore the fundamental communication costs required for achieving optimal accuracy under centralized differential privacy, Chen et al. (Chen et al., 2022a) proposed using a linear scheme based on sparse random projection to reduce communication overheads. They utilize count sketch matrices and hash functions to obtain sparse projection matrices, thus reducing model parameter dimensionality. Kerkouche et al. (Kerkouche et al., 2021) also sparsify vectors by decomposing model parameters into a “sparse orthogonal basis matrix” and a “sparse signal” multiplied by the discrete Gaussian mechanism locally. The server can reconstruct the sparse gradient vector by solving a convex quadratic optimization problem. In addition to designing new discrete noise mechanisms to integrate secure aggregation techniques, Stevens et al. (Stevens et al., 2022) proposed a secure federated aggregation scheme based on the Learning With Errors (LWE) problem (Regev, 2009). It encrypts the model parameters of local clients using masks generated by the LWE problem before uploading. This method utilizes the random noise naturally introduced by LWE technology, ensuring that the sum of errors in the LWE problem satisfies differential privacy.

### v4

@@ 2.1.2. Vertical Federated Learning (VFL)
VFL applies to scenarios where participants share the same data sample IDs but have different feature spaces. The goal is to jointly train a global model $\mathbf{w}^{\star}$ by combining features from all parties (Yang et al., 2019). Typically, one party (the server) holds the data labels, while the others (the clients) only have feature information.
(2) $\mathbf{w}^{\star}$ $\triangleq\min_{\mathbf{w}}F\left(\mathbf{w}\right),\,\text{where }\,F\left(\mathbf{w}\right)\triangleq\frac{1}{N}\sum_{i=1}^{N}L_{K}(\mathbf{w}_{K};h_{1}\left(x_{i,1},\mathbf{w}_{1}\right),\ldots,h_{K}\left(x_{i,K},\mathbf{w}_{K}\right),y_{i,K}).$
Here, $L_{K}(\cdot)$ is the loss function of the server with the label information.
@@ 2.1.3. Transfer Federated Learning (TFL)
TFL refers to FL scenarios with limited or no overlap in feature or sample spaces among participants (Liu et al., 2020c). In this setting, clients with source domain datasets and a server with a target domain dataset jointly train a model to generalize to the target domain while preserving data privacy. If the target domain lacks labels, it often requires using source domain data to generate pseudo-labels for the target domain.
@@ 2.2. DP Models
Differential privacy (DP) (Dwork et al., 2006) is a formal definition of data privacy, primarily used in centralized settings with a trusted server. It protects against privacy attacks by adding noise to the output of statistical queries. This mechanism ensures that the inclusion or exclusion of any single data record does not cause a statistically significant change in the output, thus safeguarding individual privacy.
@@ 2.2.2. LDP
From Definition 2.4, it can be seen that LDP ensures that algorithm $\mathcal{R}$ satisfies $(\epsilon,\delta)$-LDP by controlling the similarity of any two records’ output. In short, according to a certain output result of privacy algorithm $\mathcal{A}$, it is almost impossible to infer which record its input data is. In DP, its privacy algorithm $\mathcal{A}$ is defined by neighbor dataset, so it requires truthful third-party data collectors to protect data analysis results. We can see that the main difference between DP and LDP lies in the definition of neighboring datasets. LDP requires that all pairs of sensitive data must satisfy the same $\epsilon$ privacy guarantee, which may hide too much information from the datasets, leading to insufficient utility for certain applications. Gursoy et al. (Gursoy et al., 2019) extended metric-based extensions of differetial privacy to LDP and proposed condensed LDP (CLDP) to improve the utility, which measures the level of privacy guarantee between any pair of sensitive data based on their distance.
@@ 2.2.3. Shuffle Model
The shuffle differential privacy model, built upon the foundation of LDP. It adds a trusted shuffler between the server and the clients, which shuffles the data items submitted by clients to achieve anonymization. In shuffle model, each client satisfies privacy guarantee of LDP when facing shuffler, and then achieves privacy guarantee of DP when facing the server since the shuffled data forms a dataset and generates the notion of neighborhood dataset.
@@ 3.1. DP-HFL
(Client-level DP (CL-DP) [^4] Some articles treat “user-level DP” as equivalent to “client- level DP” because they assume that each client has only one user (Cheng et al., 2022; Yang et al., 2023a). However, we provide a more general definition, where we consider the possibility of one or more users under a single client. Further discussion and details regarding the definition and functionality of user-level are addressed in Section 6.). Under CL-DP, two datasets $D$ and $D^{\prime}$ are neighbouring if they differ in a single client or device (either through addition or through removal).
@@ 3.1.1. SL-DP
However, the impact of data heterogeneity is a key obstacle in DP-FL. Huang et al. (Huang et al., 2020) proposed differential privacy convolutional neural network with adaptive gradient descent algorithm (DPAGD-CNN) to update the training parameters of each client. They selected the best learning rate from a candidate set based on model evaluation in each round of local DPSGD. Noble et al. (Noble et al., 2022) introduced DP-SCAFFLOD, an extension of the SCAFFLOD algorithm (Karimireddy et al., 2020), which incorporates differential privacy. It employs a control variable to limit model drift during local DPSGD, aligning them more closely with the global model direction. Wei et al. (Wei et al., 2021) found that there is an optimal number of communication rounds in terms of convergence performance for a given privacy budget $\epsilon$, which has motivated them to adaptively allocate privacy budgets in each round. Fu et al. (Fu et al., 2022a) proposed the Adap DP-FL algorithm, which includes adaptive gradient clipping and adaptive noise scale reduction methods. In the gradient clipping step of DPSGD, gradients are clipped using adaptive thresholds to account for the heterogeneity of gradient magnitudes across clients and training rounds. Additionally, during noise addition, the noise scale gradually decreases as gradients converge across training rounds. Yang et al. (Yang et al., 2023b) start from the Non-IID data itself, by concurrently updating local data transformation layers during local model training, thereby reducing the additional heterogeneity introduced by DP, and consequently improving the utility of FL models for Non-IID data. Their method was also applied to the CL-DP setting. Chen et al. (Chen et al., 2024) utilized a adaptive server-side optimization method, simultaneously mitigating the negative impact of non-IID data and DP noise on model utility.
@@
The allocation of privacy budget is also the focus of the study, as the privacy budget determines factors such as the number of iteration rounds and noise scale, it directly impacts the performance of the trained model. Ling et al. (Ling et al., 2024) investigated how to achieve better model performance under constraints of privacy budget and communication resources. They conducted convergence analysis of DP-HFL, derived the optimal number of local iterations before each aggregation. Liu et al. (Liu et al., 2021b) considered scenarios with heterogeneous privacy budgets and proposed the Projected Federated Averaging (PFA) algorithm. This algorithm utilizes the top singular subspace of model updates submitted by clients with higher privacy budgets to project them onto model updates from clients with lower privacy budgets. Furthermore, Liu et al. (Liu et al., 2024b) pioneered the formulation of sample-level personalized differential privacy in federated learning. They introduced a two-stage hybrid sampling scheme to satisfy personalized privacy budget requirements at the sample level and provided a theoretical analysis of the privacy amplification effect for the mechanism. In addition to heterogeneous privacy budgets in each client, malekmohammadi et al. (Malekmohammadi et al., 2024) also considered the heterogeneity of client batching and dataset sizes to propose Robust-HDP, which effectively estimates the true noise level in client model updates and substantially reduces the noise in aggregated model updates. Furthermore, it reduces communication overhead by having clients with lower privacy budgets upload projected model updates instead of original model values. Zheng et al. (Zheng et al., 2021) integrated GDP privacy metric into DP-HFL, proposing a private federated learning framework called PriFedSync. While considering the cost of communication, Li et al. (Li et al., 2022c) proposed SoteriaFL, which is a unified framework for compressed private FL. They use shifted compression scheme (Horváth et al., 2023) to compress the perturbed parameters for efficient communication while maintaining high utility.
@@ 3.1.2. CL-DP
In addition to this, many studies have focused on reducing the impact of noise on uploaded model parameters. Chen et al. (Cheng et al., 2022) observed that using a small clipping threshold can decrease the injected noise volume. They reduced the norm of local updates by regularizing local models and making the local updates sparse. DP-FedSAM  (Shi et al., 2023) uses a SAM optimizer (Foret et al., 2020) to help model parameters escape saddle points and enhance the robustness of the model to noise, aiming to find more stable convergence points. Bietti et al. (Bietti et al., 2022) proposed PPSGD, which trains personalized local models and ensures global model privacy simultaneously. They utilized personalized models to enhance the global model’s performance. Similarly, Yang et al. (Yang et al., 2023a) dynamically preserved high-information model parameters locally from noise impact using layer-wise Fisher information. They also introduced an adaptive regularization strategy to impose differential constraints on model parameters uploaded to the server, enhancing robustness to clipping. Xu et al. (Xu et al., 2023a) took into account that the size of the softmax layer of model parameters is linearly related to the number of sample labels. By keeping the softmax layer local and not uploading it, the participation of more clients (i.e., more sample labels) in federated learning does not result in more noise injection. Triastcyn et al. (Triastcyn and Faltings, 2019) leveraged the assumption that data is distributed similarly across clients in HFL, which leads to their updates being highly consistent, and utilized BDP for privacy auditing to obtain tighter privacy bounds. Their assumption and method is not only used under the CL-DP, but also applied in SL-DP scenario.
CL-DP with Secure Aggregation (SA)  [^7] Many works refer to this framework as “distributed dp” (Kairouz et al., 2021a; Agarwal et al., 2021; Yang et al., 2023b).. As mentioned above, CL-DP often requires a secure server for aggregation. In order to eliminate the reliance on a trusted central authority, many recent studies have combined differential privacy techniques with secure aggregation (Bonawitz et al., 2017) to achieve CL-DP. As shown in Figure 4, after local training, each client adds a small amount of differential privacy noise to the model parameters, then encrypts and sends them to the server. The server performs aggregation, enabling the aggregation of noise levels from each client, thereby ensuring that the encrypted model satisfies CL-DP. However, secure aggregation relies on modular arithmetic and cannot be directly compatible with continuous noise mechanisms (such as Laplace mechanism and Gaussian mechanism) currently used to implement differential privacy. This has led researchers to begin designing discrete noise addition mechanisms and integrating them with secure aggregation to characterize differential privacy.
@@
However, the above methods rely on per-parameter quantization, resulting in significant communication overheads. To further explore the fundamental communication costs required for achieving optimal accuracy under centralized differential privacy, Chen et al. (Chen et al., 2022a) proposed using a linear scheme based on sparse random projection to reduce communication overheads. They utilize count sketch matrices and hash functions to obtain sparse projection matrices, thus reducing model parameter dimensionality. Kerkouche et al. (Kerkouche et al., 2021) also sparsify vectors by decomposing model parameters into a “sparse orthogonal basis matrix” and a “sparse signal” multiplied by the discrete Gaussian mechanism locally. The server can reconstruct the sparse gradient vector by solving a convex quadratic optimization problem. In addition to designing new discrete noise mechanisms to integrate secure aggregation techniques, Stevens et al. (Stevens et al., 2022) proposed a secure federated aggregation scheme based on the Learning With Errors (LWE) problem (Regev, 2009). It encrypts the model parameters of local clients using masks generated by the LWE problem before uploading. This method utilizes the random noise naturally introduced by LWE technology, ensuring that the sum of errors in the LWE problem satisfies differential privacy. Kato et al. (Kato et al., 2023) implement CL-DP by constructing a trusted aggregator using Trusted Execution Environment (TEE).